# Structure-Preserving Oscillation-Eliminating Discontinuous Galerkin Schemes for Ideal MHD Equations: Locally Divergence-Free and Positivity-Preserving

### v1

@@ 3.6 Rigorous PP analysis
In order to prove that $\bar{\boldsymbol{U}}_{ij}^{n+1}\in\mathcal{G}$ in scheme (55), we first provide several key auxiliary theories, and the rigorous proof of the PP property of the scheme will be given by Theorem 1 later.
Lemma 1.
The admissible state set $\mathcal{G}$ is equivalent to
$\mathcal{G}_{*}=\left\{\boldsymbol{U}=\left(\rho,\boldsymbol{m},\boldsymbol{B},E\right)^{\top}:\rho>0,\boldsymbol{U}\cdot\boldsymbol{n}^{*}+\frac{\left|\boldsymbol{B}^{*}\right|^{2}}{2}>0,\forall\boldsymbol{v}^{*},\boldsymbol{B}^{*}\in\mathbb{R}^{3}\right\},$
where
$\boldsymbol{n}^{*}=\left(\frac{\left|\boldsymbol{v}^{*}\right|^{2}}{2},-\boldsymbol{v}^{*},-\boldsymbol{B}^{*},1\right)^{\top}.$
The proof of Lemma 1 can be found in [42]. For more details about its geometric interpretation, please see the geometric quasilinearization (GQL) approach in [45].
Lemma 2.
@@
$\left|\sqrt{\rho}\left(\boldsymbol{v}-\boldsymbol{v}^{*}\right)\cdot\left(\boldsymbol{B}-\boldsymbol{B}^{*}\right)\right|<\boldsymbol{U}\cdot\boldsymbol{n}^{*}+\frac{\left|\boldsymbol{B}^{*}\right|^{2}}{2}.$
@@
$-b\left(\boldsymbol{S}\left(\boldsymbol{U}\right)\cdot\boldsymbol{n}^{*}\right)\geq b\left(\boldsymbol{v}^{*}\cdot\boldsymbol{B}^{*}\right)-\frac{|b|}{\sqrt{\rho}}\left(\boldsymbol{U}\cdot\boldsymbol{n}^{*}+\frac{\left|\boldsymbol{B}^{*}\right|^{2}}{2}\right).$ (56)
@@
Corollary 1.
Given admissible states $\boldsymbol{U}^{(j)}$, $1\leq j\leq 4$, we define
$\hat{\alpha}_{1}:=\mathscr{C}_{1}\left(\boldsymbol{U}^{(1)}\right)+\frac{\Delta y}{\Delta x+\Delta y}\frac{\left|\boldsymbol{B}^{(1)}-\boldsymbol{B}^{(2)}\right|}{\sqrt{\rho^{(1)}}+\sqrt{\rho^{(2)}}}+\frac{\Delta x}{\Delta x+\Delta y}\frac{\left|\boldsymbol{B}^{(1)}-\boldsymbol{B}^{(3)}\right|}{\sqrt{\rho^{(1)}}+\sqrt{\rho^{(3)}}}+\frac{\Delta x}{\Delta x+\Delta y}\frac{\left|\boldsymbol{B}^{(1)}-\boldsymbol{B}^{(4)}\right|}{\sqrt{\rho^{(1)}}+\sqrt{\rho^{(4)}}}+$
$+\max\left\{\left(u_{1}\right)^{(1)},\frac{\Delta y}{\Delta x+\Delta y}\frac{\sqrt{\rho^{(1)}}\left(u_{1}\right)^{(1)}+\sqrt{\rho^{(2)}}\left(u_{1}\right)^{(2)}}{\sqrt{\rho^{(1)}}+\sqrt{\rho^{(2)}}}\right.$
$+\frac{\Delta x}{2\left(\Delta x+\Delta y\right)}\frac{\sqrt{\rho^{(1)}}\left(\left(u_{1}\right)^{(1)}-\left(u_{2}\right)^{(1)}\right)+\sqrt{\rho^{(3)}}\left(\left(u_{1}\right)^{(3)}-\left(u_{2}\right)^{(3)}\right)}{\sqrt{\rho^{(1)}}+\sqrt{\rho^{(3)}}}$
$+\left.\frac{\Delta x}{2\left(\Delta x+\Delta y\right)}\frac{\sqrt{\rho^{(1)}}\left(\left(u_{1}\right)^{(1)}+\left(u_{2}\right)^{(1)}\right)+\sqrt{\rho^{(4)}}\left(\left(u_{1}\right)^{(4)}+\left(u_{2}\right)^{(4)}\right)}{\sqrt{\rho^{(1)}}+\sqrt{\rho^{(4)}}}\right\},$
$\hat{\alpha}_{2}:=\mathscr{C}_{1}\left(\boldsymbol{U}^{(2)}\right)+\frac{\Delta y}{\Delta x+\Delta y}\frac{\left|\boldsymbol{B}^{(2)}-\boldsymbol{B}^{(1)}\right|}{\sqrt{\rho^{(2)}}+\sqrt{\rho^{(1)}}}+\frac{\Delta x}{\Delta x+\Delta y}\frac{\left|\boldsymbol{B}^{(2)}-\boldsymbol{B}^{(3)}\right|}{\sqrt{\rho^{(2)}}+\sqrt{\rho^{(3)}}}+\frac{\Delta x}{\Delta x+\Delta y}\frac{\left|\boldsymbol{B}^{(2)}-\boldsymbol{B}^{(4)}\right|}{\sqrt{\rho^{(2)}}+\sqrt{\rho^{(4)}}}+$
$+\max\left\{-\left(u_{1}\right)^{(2)},\frac{-\Delta y}{\Delta x+\Delta y}\frac{\sqrt{\rho^{(2)}}\left(u_{1}\right)^{(2)}+\sqrt{\rho^{(1)}}\left(u_{1}\right)^{(1)}}{\sqrt{\rho^{(2)}}+\sqrt{\rho^{(1)}}}\right.$
$+\frac{-\Delta x}{2\left(\Delta x+\Delta y\right)}\frac{\sqrt{\rho^{(2)}}\left(\left(u_{1}\right)^{(2)}+\left(u_{2}\right)^{(2)}\right)+\sqrt{\rho^{(3)}}\left(\left(u_{1}\right)^{(3)}+\left(u_{2}\right)^{(3)}\right)}{\sqrt{\rho^{(2)}}+\sqrt{\rho^{(3)}}}$
$+\left.\frac{\Delta x}{2\left(\Delta x+\Delta y\right)}\frac{\sqrt{\rho^{(2)}}\left(\left(u_{2}\right)^{(2)}-\left(u_{1}\right)^{(2)}\right)+\sqrt{\rho^{(4)}}\left(\left(u_{2}\right)^{(4)}-\left(u_{1}\right)^{(4)}\right)}{\sqrt{\rho^{(2)}}+\sqrt{\rho^{(4)}}}\right\},$
$\hat{\alpha}_{3}:=\mathscr{C}_{2}\left(\boldsymbol{U}^{(3)}\right)+\frac{\Delta x}{\Delta x+\Delta y}\frac{\left|\boldsymbol{B}^{(3)}-\boldsymbol{B}^{(4)}\right|}{\sqrt{\rho^{(3)}}+\sqrt{\rho^{(4)}}}+\frac{\Delta y}{\Delta x+\Delta y}\frac{\left|\boldsymbol{B}^{(3)}-\boldsymbol{B}^{(1)}\right|}{\sqrt{\rho^{(3)}}+\sqrt{\rho^{(1)}}}+\frac{\Delta y}{\Delta x+\Delta y}\frac{\left|\boldsymbol{B}^{(3)}-\boldsymbol{B}^{(2)}\right|}{\sqrt{\rho^{(3)}}+\sqrt{\rho^{(2)}}}+$
$+\max\left\{\left(u_{2}\right)^{(3)},\frac{\Delta x}{\Delta x+\Delta y}\frac{\sqrt{\rho^{(3)}}\left(u_{2}\right)^{(3)}+\sqrt{\rho^{(4)}}\left(u_{2}\right)^{(4)}}{\sqrt{\rho^{(3)}}+\sqrt{\rho^{(4)}}}\right.$
$+\frac{\Delta y}{2\left(\Delta x+\Delta y\right)}\frac{\sqrt{\rho^{(3)}}\left(-\left(u_{1}\right)^{(3)}+\left(u_{2}\right)^{(3)}\right)+\sqrt{\rho^{(1)}}\left(-\left(u_{1}\right)^{(1)}+\left(u_{2}\right)^{(1)}\right)}{\sqrt{\rho^{(3)}}+\sqrt{\rho^{(1)}}}$
$+\left.\frac{\Delta y}{2\left(\Delta x+\Delta y\right)}\frac{\sqrt{\rho^{(3)}}\left(\left(u_{1}\right)^{(3)}+\left(u_{2}\right)^{(3)}\right)+\sqrt{\rho^{(2)}}\left(\left(u_{1}\right)^{(2)}+\left(u_{2}\right)^{(2)}\right)}{\sqrt{\rho^{(3)}}+\sqrt{\rho^{(2)}}}\right\},$
$\hat{\alpha}_{4}:=\mathscr{C}_{2}\left(\boldsymbol{U}^{(4)}\right)+\frac{\Delta x}{\Delta x+\Delta y}\frac{\left|\boldsymbol{B}^{(4)}-\boldsymbol{B}^{(3)}\right|}{\sqrt{\rho^{(4)}}+\sqrt{\rho^{(3)}}}+\frac{\Delta y}{\Delta x+\Delta y}\frac{\left|\boldsymbol{B}^{(4)}-\boldsymbol{B}^{(1)}\right|}{\sqrt{\rho^{(4)}}+\sqrt{\rho^{(1)}}}+\frac{\Delta y}{\Delta x+\Delta y}\frac{\left|\boldsymbol{B}^{(4)}-\boldsymbol{B}^{(2)}\right|}{\sqrt{\rho^{(4)}}+\sqrt{\rho^{(2)}}}+$
$+\max\left\{-\left(u_{2}\right)^{(4)},\frac{-\Delta x}{\Delta x+\Delta y}\frac{\sqrt{\rho^{(4)}}\left(u_{2}\right)^{(4)}+\sqrt{\rho^{(3)}}\left(u_{2}\right)^{(3)}}{\sqrt{\rho^{(4)}}+\sqrt{\rho^{(3)}}}\right.$
$+\frac{-\Delta y}{2\left(\Delta x+\Delta y\right)}\frac{\sqrt{\rho^{(4)}}\left(\left(u_{1}\right)^{(4)}+\left(u_{2}\right)^{(4)}\right)+\sqrt{\rho^{(1)}}\left(\left(u_{1}\right)^{(1)}+\left(u_{2}\right)^{(1)}\right)}{\sqrt{\rho^{(4)}}+\sqrt{\rho^{(1)}}}$
$+\left.\frac{\Delta y}{2\left(\Delta x+\Delta y\right)}\frac{\sqrt{\rho^{(4)}}\left(\left(u_{1}\right)^{(4)}-\left(u_{2}\right)^{(4)}\right)+\sqrt{\rho^{(2)}}\left(\left(u_{1}\right)^{(2)}-\left(u_{2}\right)^{(2)}\right)}{\sqrt{\rho^{(4)}}+\sqrt{\rho^{(2)}}}\right\}.$
Then for any $\alpha_{j}\geq\hat{\alpha}_{j}$, the state
$\bar{\boldsymbol{U}}:=\frac{1}{\Delta y\left(\alpha_{1}+\alpha_{2}\right)+\Delta x\left(\alpha_{3}+\alpha_{4}\right)}\left[\Delta y\left(\alpha_{1}\boldsymbol{U}^{(1)}-\boldsymbol{F}_{1}\left(\boldsymbol{U}^{(1)}\right)\right)+\Delta y\left(\alpha_{2}\boldsymbol{U}^{(2)}+\boldsymbol{F}_{1}\left(\boldsymbol{U}^{(2)}\right)\right)\right.$
$+\left.\Delta x\left(\alpha_{3}\boldsymbol{U}^{(3)}-\boldsymbol{F}_{2}\left(\boldsymbol{U}^{(3)}\right)\right)+\Delta x\left(\alpha_{4}\boldsymbol{U}^{(4)}+\boldsymbol{F}_{2}\left(\boldsymbol{U}^{(4)}\right)\right)\right],$
belongs to $\mathcal{G}_{\rho}:=\left\{\boldsymbol{U}=\left(\rho,\boldsymbol{m},\boldsymbol{B},E\right)^{\top}:\rho>0\right\}$, and satisfies
@@
$\bar{\boldsymbol{U}}\cdot\boldsymbol{n}^{*}+\frac{\left|\boldsymbol{B}^{*}\right|^{2}}{2}\geq\frac{-\boldsymbol{v}^{*}\cdot\boldsymbol{B}^{*}\left(\Delta y\left(B_{1}\right)^{(1)}-\Delta y\left(B_{1}\right)^{(2)}+\Delta x\left(B_{2}\right)^{(3)}-\Delta x\left(B_{2}\right)^{(4)}\right)}{\Delta y\left(\alpha_{1}+\alpha_{2}\right)+\Delta x\left(\alpha_{3}+\alpha_{4}\right)},\forall\boldsymbol{v}^{*},\boldsymbol{B}^{*}\in\mathbb{R}^{3}.$
Furthermore, $\Delta y\left(\alpha_{1}+\alpha_{2}\right)+\Delta x\left(\alpha_{3}+\alpha_{4}\right)>0$.
@@
Proof This directly follows from Theorem 1 in [44] with $N=4$, by taking
$s_{1}=s_{2}=\Delta y,\quad s_{3}=s_{4}=\Delta x,\quad\boldsymbol{\xi}^{(1)}=(1,0),\quad\boldsymbol{\xi}^{(2)}=(-1,0),\quad\boldsymbol{\xi}^{(3)}=(0,1),\quad\boldsymbol{\xi}^{(4)}=(0,-1).\qed$
Corollary 2.
@@
$\mathscr{V}_{\ell,r}\geq\alpha_{r}\left(\boldsymbol{U}^{+},\boldsymbol{U}^{-}\right),\quad\mathscr{V}_{\ell,l}\geq\alpha_{l}\left(\boldsymbol{U}^{-},\boldsymbol{U}^{+}\right),\quad\ell=1,2,$ (57)
@@
$\widehat{\boldsymbol{F}_{\ell}}\left(\boldsymbol{U}^{-},\boldsymbol{U}^{+}\right)=\mathscr{V}_{\ell}^{-}\boldsymbol{H}_{\ell}\left(\boldsymbol{U}^{-},\boldsymbol{U}^{+}\right)+\boldsymbol{F}_{\ell}\left(\boldsymbol{U}^{-}\right)-\mathscr{V}_{\ell}^{-}\boldsymbol{U}^{-}$ (58)
@@
where the intermediate state
@@
belongs to $\mathcal{G}$ and satisfies
$\boldsymbol{H}_{\ell}\left(\boldsymbol{U}^{-},\boldsymbol{U}^{+}\right)\cdot\boldsymbol{n}^{*}+\frac{\left|\boldsymbol{B}^{*}\right|^{2}}{2}\geq-\frac{\boldsymbol{v}^{*}\cdot\boldsymbol{B}^{*}}{\mathscr{V}_{\ell}^{+}-\mathscr{V}_{\ell}^{-}}\left(\boldsymbol{B}_{\ell}^{+}-\boldsymbol{B}_{\ell}^{-}\right),\quad\forall\boldsymbol{v}^{*},\boldsymbol{B}^{*}\in\mathbb{R}^{3}.$ (60)
@@
With the above auxiliary theories as the key cornerstone, we now proof the PP property of the scheme (55). To shorten the notations, we define
@@
$\hat{\alpha}_{1,i\pm\frac{1}{2}}^{\mu,\mp}=\mathscr{C}_{1}\left(\boldsymbol{U}_{i\pm\frac{1}{2},j}^{\mp,\mu}\right)+\frac{\Delta y}{\Delta x+\Delta y}\frac{\left|\boldsymbol{B}_{i\pm\frac{1}{2},j}^{\mp,\mu}-\boldsymbol{B}_{i\mp\frac{1}{2},j}^{\pm,\mu}\right|}{\sqrt{\rho_{i\pm\frac{1}{2},j}^{\mp,\mu}}+\sqrt{\rho_{i\mp\frac{1}{2},j}^{\pm,\mu}}}+\frac{\Delta x}{\Delta x+\Delta y}\frac{\left|\boldsymbol{B}_{i\pm\frac{1}{2},j}^{\mp,\mu}-\boldsymbol{B}_{i,j+\frac{1}{2}}^{\mu,-}\right|}{\sqrt{\rho_{i\pm\frac{1}{2},j}^{\mp,\mu}}+\sqrt{\rho_{i,j+\frac{1}{2}}^{\mu,-}}}$
$+\frac{\Delta x}{\Delta x+\Delta y}\frac{\left|\boldsymbol{B}_{i\pm\frac{1}{2},j}^{\mp,\mu}-\boldsymbol{B}_{i,j-\frac{1}{2}}^{\mu,+}\right|}{\sqrt{\rho_{i\pm\frac{1}{2},j}^{\mp,\mu}}+\sqrt{\rho_{i,j-\frac{1}{2}}^{\mu,+}}}$
$+\max\left\{\pm\left(u_{1}\right)_{i\pm\frac{1}{2},j}^{\mp,\mu},\frac{\pm\Delta y}{\left(\Delta x+\Delta y\right)}\frac{\sqrt{\rho_{i\pm\frac{1}{2},j}^{\mp,\mu}}\left(u_{1}\right)_{i\pm\frac{1}{2},j}^{\mp,\mu}+\sqrt{\rho_{i\mp\frac{1}{2},j}^{\pm,\mu}}\left(u_{1}\right)_{i\mp\frac{1}{2},j}^{\pm,\mu}}{\sqrt{\rho_{i\pm\frac{1}{2},j}^{\mp,\mu}}+\sqrt{\rho_{i\mp\frac{1}{2},j}^{\pm,\mu}}}\right.$
$+\frac{\Delta x}{2\left(\Delta x+\Delta y\right)}\frac{\sqrt{\rho_{i\pm\frac{1}{2},j}^{\mp,\mu}}\left(\pm\left(u_{1}\right)_{i\pm\frac{1}{2},j}^{\mp,\mu}-\left(u_{2}\right)_{i\pm\frac{1}{2},j}^{\mp,\mu}\right)+\sqrt{\rho_{i,j+\frac{1}{2}}^{\mu,-}}\left(\pm\left(u_{1}\right)_{i,j+\frac{1}{2}}^{\mu,-}-\left(u_{2}\right)_{i,j+\frac{1}{2}}^{\mu,-}\right)}{\sqrt{\rho_{i\pm\frac{1}{2},j}^{\mp,\mu}}+\sqrt{\rho_{i,j+\frac{1}{2}}^{\mu,-}}}$
$+\left.\frac{\Delta x}{2\left(\Delta x+\Delta y\right)}\frac{\sqrt{\rho_{i\pm\frac{1}{2},j}^{\mp,\mu}}\left(\pm\left(u_{1}\right)_{i\pm\frac{1}{2},j}^{\mp,\mu}+\left(u_{2}\right)_{i\pm\frac{1}{2},j}^{\mp,\mu}\right)+\sqrt{\rho_{i,j-\frac{1}{2}}^{\mu,+}}\left(\pm\left(u_{1}\right)_{i,j-\frac{1}{2}}^{\mu,+}+\left(u_{2}\right)_{i,j-\frac{1}{2}}^{\mu,+}\right)}{\sqrt{\rho_{i\pm\frac{1}{2},j}^{\mp,\mu}}+\sqrt{\rho_{i,j-\frac{1}{2}}^{\mu,+}}}\right\},$
$\hat{\alpha}_{2,j\pm\frac{1}{2}}^{\mu,\mp}=\mathscr{C}_{2}\left(\boldsymbol{U}_{i,j\pm\frac{1}{2}}^{\mu,\mp}\right)+\frac{\Delta x}{\Delta x+\Delta y}\frac{\left|\boldsymbol{B}_{i,j\pm\frac{1}{2}}^{\mu,\mp}-\boldsymbol{B}_{i,j\mp\frac{1}{2}}^{\mu,\pm}\right|}{\sqrt{\rho_{i,j\pm\frac{1}{2}}^{\mu,\mp}}+\sqrt{\rho_{i,j\mp\frac{1}{2}}^{\mu,\pm}}}+\frac{\Delta y}{\Delta x+\Delta y}\frac{\left|\boldsymbol{B}_{i,j\pm\frac{1}{2}}^{\mu,\mp}-\boldsymbol{B}_{i+\frac{1}{2},j}^{-,\mu}\right|}{\sqrt{\rho_{i,j\pm\frac{1}{2}}^{\mu,\mp}}+\sqrt{\rho_{i+\frac{1}{2},j}^{-,\mu}}}$
$+\frac{\Delta y}{\Delta x+\Delta y}\frac{\left|\boldsymbol{B}_{i,j\pm\frac{1}{2}}^{\mu,\mp}-\boldsymbol{B}_{i-\frac{1}{2},j}^{+,\mu}\right|}{\sqrt{\rho_{i,j\pm\frac{1}{2}}^{\mu,\mp}}+\sqrt{\rho_{i-\frac{1}{2},j}^{+,\mu}}}$
$+\max\left\{\pm\left(u_{2}\right)_{i,j\pm\frac{1}{2}}^{\mu,\mp},\frac{\pm\Delta x}{\left(\Delta x+\Delta y\right)}\frac{\sqrt{\rho_{i,j\pm\frac{1}{2}}^{\mu,\mp}}\left(u_{2}\right)_{i,j\pm\frac{1}{2}}^{\mu,\mp}+\sqrt{\rho_{i,j\mp\frac{1}{2}}^{\mu,\pm}}\left(u_{2}\right)_{i,j\mp\frac{1}{2}}^{\mu,\pm}}{\sqrt{\rho_{i,j\pm\frac{1}{2}}^{\mu,\mp}}+\sqrt{\rho_{i,j\mp\frac{1}{2}}^{\mu,\pm}}}\right.$
$+\frac{\Delta y}{2\left(\Delta x+\Delta y\right)}\frac{\sqrt{\rho_{i,j\pm\frac{1}{2}}^{\mu,\mp}}\left(-\left(u_{1}\right)_{i,j\pm\frac{1}{2}}^{\mu,\mp}\pm\left(u_{2}\right)_{i,j\pm\frac{1}{2}}^{\mu,\mp}\right)+\sqrt{\rho_{i+\frac{1}{2},j}^{-,\mu}}\left(-\left(u_{1}\right)_{i+\frac{1}{2},j}^{-,\mu}\pm\left(u_{2}\right)_{i+\frac{1}{2},j}^{-,\mu}\right)}{\sqrt{\rho_{i,j\pm\frac{1}{2}}^{\mu,\mp}}+\sqrt{\rho_{i+\frac{1}{2},j}^{-,\mu}}}$
$+\left.\frac{\Delta y}{2\left(\Delta x+\Delta y\right)}\frac{\sqrt{\rho_{i,j\pm\frac{1}{2}}^{\mu,\mp}}\left(\left(u_{1}\right)_{i,j\pm\frac{1}{2}}^{\mu,\mp}\pm\left(u_{2}\right)_{i,j\pm\frac{1}{2}}^{\mu,\mp}\right)+\sqrt{\rho_{i-\frac{1}{2},j}^{+,\mu}}\left(\left(u_{1}\right)_{i-\frac{1}{2},j}^{+,\mu}\pm\left(u_{2}\right)_{i-\frac{1}{2},j}^{+,\mu}\right)}{\sqrt{\rho_{i,j\pm\frac{1}{2}}^{\mu,\mp}}+\sqrt{\rho_{i-\frac{1}{2},j}^{+,\mu}}}\right\}.$
The PP property of the scheme (55) is shown as follows.

### v2

@@ 3.6 Rigorous PP analysis
To prove $\bar{\boldsymbol{U}}_{ij}^{n+1}\in\mathcal{G}$ in scheme (59), we first introduce several key auxiliary lemmas related to the GQL approach [46, 49], which will be used in the proof of $\bar{\boldsymbol{U}}_{ij}^{n+1}\in\mathcal{G}$ in Theorem 1.
Lemma 1 (GQL representation [46]).
The admissible state set $\mathcal{G}$ is exactly equivalent to
$\mathcal{G}_{*}=\left\{\boldsymbol{U}=\left(\rho,\boldsymbol{m},\boldsymbol{B},E\right)^{\top}:~{}\rho>0,~{}\boldsymbol{U}\cdot\boldsymbol{n}^{*}+\frac{\left\|\boldsymbol{B}^{*}\right\|^{2}}{2}>0~{}~{}\forall\boldsymbol{v}^{*},\boldsymbol{B}^{*}\in\mathbb{R}^{3}\right\},$ (60)
where $\boldsymbol{v}^{*}$ and $\boldsymbol{B}^{*}$ are the extra free auxiliary variables independent of $\boldsymbol{U}$, and
$\boldsymbol{n}^{*}=\left(\frac{\left\|\boldsymbol{v}^{*}\right\|^{2}}{2},~{}-\boldsymbol{v}^{*},~{}-\boldsymbol{B}^{*},~{}1\right)^{\top}.$
In the original definition (36) of $\mathcal{G}$, the second constraint is nonlinear respect to $\boldsymbol{U}$, leading to challenges in analyzing the positivity of $\mathcal{E}(\boldsymbol{U})$. Thanks to the GQL representation (60), all equivalent constraints in $\mathcal{G}_{*}$ become linear with respect to $\boldsymbol{U}$, thereby greatly facilitating the PP analysis. The first proof of Lemma 1 can be found in [46]. For more details about its geometric interpretation, refer to the GQL framework in [49].
Lemma 2 ([47]).
@@
$\left|\sqrt{\rho}\left(\boldsymbol{v}-\boldsymbol{v}^{*}\right)\cdot\left(\boldsymbol{B}-\boldsymbol{B}^{*}\right)\right|<\boldsymbol{U}\cdot\boldsymbol{n}^{*}+\frac{\left\|\boldsymbol{B}^{*}\right\|^{2}}{2}.$
@@
$-b\left(\boldsymbol{S}\left(\boldsymbol{U}\right)\cdot\boldsymbol{n}^{*}\right)\geq b\left(\boldsymbol{v}^{*}\cdot\boldsymbol{B}^{*}\right)-\frac{|b|}{\sqrt{\rho}}\left(\boldsymbol{U}\cdot\boldsymbol{n}^{*}+\frac{\left\|\boldsymbol{B}^{*}\right\|^{2}}{2}\right).$ (61)
@@
Lemma 3.
Given two admissible states $\boldsymbol{U}^{(1)}$ and $\boldsymbol{U}^{(2)}$, we define for $\ell\in\{1,2\}$ that
$\hat{\alpha}_{\ell}^{(1)}:=\max\left\{\left(u_{\ell}\right)^{(1)},\frac{\sqrt{\rho^{(1)}}\left(u_{\ell}\right)^{(1)}+\sqrt{\rho^{(2)}}\left(u_{\ell}\right)^{(2)}}{\sqrt{\rho^{(1)}}+\sqrt{\rho^{(2)}}}\right\}+\mathscr{C}_{\ell}\left(\boldsymbol{U}^{(1)}\right)+\frac{\left\|\boldsymbol{B}^{(1)}-\boldsymbol{B}^{(2)}\right\|}{\sqrt{\rho^{(1)}}+\sqrt{\rho^{(2)}}},$
$\hat{\alpha}_{\ell}^{(2)}:=\max\left\{-\left(u_{\ell}\right)^{(2)},-\left(\frac{\sqrt{\rho^{(2)}}\left(u_{\ell}\right)^{(2)}+\sqrt{\rho^{(1)}}\left(u_{\ell}\right)^{(1)}}{\sqrt{\rho^{(2)}}+\sqrt{\rho^{(1)}}}\right)\right\}+\mathscr{C}_{\ell}\left(\boldsymbol{U}^{(2)}\right)+\frac{\left\|\boldsymbol{B}^{(2)}-\boldsymbol{B}^{(1)}\right\|}{\sqrt{\rho^{(2)}}+\sqrt{\rho^{(1)}}}.$
Then for any $\alpha_{\ell}^{(j)}\geq\hat{\alpha}_{\ell}^{(j)}$, $j=1,2$, the states
$\bar{\boldsymbol{U}}_{\ell}:=\frac{1}{\alpha_{\ell}^{(1)}+\alpha_{\ell}^{(2)}}\left[\left(\alpha_{\ell}^{(1)}\boldsymbol{U}^{(1)}-\boldsymbol{F}_{\ell}\left(\boldsymbol{U}^{(1)}\right)\right)+\left(\alpha_{\ell}^{(2)}\boldsymbol{U}^{(2)}+\boldsymbol{F}_{\ell}\left(\boldsymbol{U}^{(2)}\right)\right)\right],\quad\ell=1,2,$
belong to $\mathcal{G}_{\rho}:=\left\{\boldsymbol{U}=\left(\rho,\boldsymbol{m},\boldsymbol{B},E\right)^{\top}:\rho>0\right\}$, and satisfy
@@
$\bar{\boldsymbol{U}}_{\ell}\cdot\boldsymbol{n}^{*}+\frac{\left|\boldsymbol{B}^{*}\right|^{2}}{2}\geq\frac{-\boldsymbol{v}^{*}\cdot\boldsymbol{B}^{*}}{\alpha_{\ell}^{(1)}+\alpha_{\ell}^{(2)}}\left(\left(B_{\ell}\right)^{(1)}-\left(B_{\ell}\right)^{(2)}\right)\quad\forall\boldsymbol{v}^{*},\boldsymbol{B}^{*}\in\mathbb{R}^{3}.$
@@
Proof This directly follows from Theorem 1 in [48] with $N=2$, by taking $s_{1}=s_{2}=1$, $\boldsymbol{\xi}^{(1)}=-\boldsymbol{\xi}^{(2)}=(1,0)$ for $\ell=1$, and $\boldsymbol{\xi}^{(1)}=-\boldsymbol{\xi}^{(2)}=(0,1)$ for $\ell=2.$
Lemma 4.
@@
$\mathscr{V}_{\ell,r}\geq\alpha_{\ell,r}\left(\boldsymbol{U}^{+},\boldsymbol{U}^{-}\right),\quad\mathscr{V}_{\ell,l}\leq\alpha_{\ell,l}\left(\boldsymbol{U}^{-},\boldsymbol{U}^{+}\right),\quad\ell=1,2,$ (62)
@@
$\widehat{\boldsymbol{F}_{\ell}}\left(\boldsymbol{U}^{-},\boldsymbol{U}^{+}\right)=\mathscr{V}_{\ell}^{-}\boldsymbol{H}_{\ell}\left(\boldsymbol{U}^{-},\boldsymbol{U}^{+}\right)+\boldsymbol{F}_{\ell}\left(\boldsymbol{U}^{-}\right)-\mathscr{V}_{\ell}^{-}\boldsymbol{U}^{-},$ (63)
@@
with the intermediate state
@@
belonging to $\mathcal{G}$ and satisfying
$\boldsymbol{H}_{\ell}\left(\boldsymbol{U}^{-},\boldsymbol{U}^{+}\right)\cdot\boldsymbol{n}^{*}+\frac{\left\|\boldsymbol{B}^{*}\right\|^{2}}{2}\geq-\frac{\boldsymbol{v}^{*}\cdot\boldsymbol{B}^{*}}{\mathscr{V}_{\ell}^{+}-\mathscr{V}_{\ell}^{-}}\left(\boldsymbol{B}_{\ell}^{+}-\boldsymbol{B}_{\ell}^{-}\right)\quad\forall\boldsymbol{v}^{*},\boldsymbol{B}^{*}\in\mathbb{R}^{3}.$ (65)
@@
With the above auxiliary lemmas as the key cornerstone, we are now ready to prove $\bar{\boldsymbol{U}}_{ij}^{n+1}\in\mathcal{G}$ in scheme (59). Recall that $\widetilde{\boldsymbol{U}}_{\sigma}^{n}$ denotes the OEDG solution after the LDF OE and PP limiting procedures. To shorten our notations, define
@@
$\hat{\alpha}_{1,i\pm\frac{1}{2}}^{\mu,\mp}=$ $\max\left\{\pm\left(u_{1}\right)_{i\pm\frac{1}{2},j}^{\mp,\mu},\pm\left(\frac{\sqrt{\rho_{i\pm\frac{1}{2},j}^{\mp,\mu}}\left(u_{1}\right)_{i\pm\frac{1}{2},j}^{\mp,\mu}+\sqrt{\rho_{i\mp\frac{1}{2},j}^{\pm,\mu}}\left(u_{1}\right)_{i\mp\frac{1}{2},j}^{\pm,\mu}}{\sqrt{\rho_{i\pm\frac{1}{2},j}^{\mp,\mu}}+\sqrt{\rho_{i\mp\frac{1}{2},j}^{\pm,\mu}}}\right)\right\}$
$+\mathscr{C}_{1}\left(\boldsymbol{U}_{i\pm\frac{1}{2},j}^{\mp,\mu}\right)+\frac{\left\|\boldsymbol{B}_{i\pm\frac{1}{2},j}^{\mp,\mu}-\boldsymbol{B}_{i\mp\frac{1}{2},j}^{\pm,\mu}\right\|}{\sqrt{\rho_{i\pm\frac{1}{2},j}^{\mp,\mu}}+\sqrt{\rho_{i\mp\frac{1}{2},j}^{\pm,\mu}}},$
$\hat{\alpha}_{2,j\pm\frac{1}{2}}^{\mu,\mp}=$ $\max\left\{\pm\left(u_{2}\right)_{i,j\pm\frac{1}{2}}^{\mu,\mp},\pm\left(\frac{\sqrt{\rho_{i,j\pm\frac{1}{2}}^{\mu,\mp}}\left(u_{2}\right)_{i,j\pm\frac{1}{2}}^{\mu,\mp}+\sqrt{\rho_{i,j\mp\frac{1}{2}}^{\mu,\pm}}\left(u_{2}\right)_{i,j\mp\frac{1}{2}}^{\mu,\pm}}{\sqrt{\rho_{i,j\pm\frac{1}{2}}^{\mu,\mp}}+\sqrt{\rho_{i,j\mp\frac{1}{2}}^{\mu,\pm}}}\right)\right\}$
$+\mathscr{C}_{2}\left(\boldsymbol{U}_{i,j\pm\frac{1}{2}}^{\mu,\mp}\right)+\frac{\left\|\boldsymbol{B}_{i,j\pm\frac{1}{2}}^{\mu,\mp}-\boldsymbol{B}_{i,j\mp\frac{1}{2}}^{\mu,\pm}\right\|}{\sqrt{\rho_{i,j\pm\frac{1}{2}}^{\mu,\mp}}+\sqrt{\rho_{i,j\mp\frac{1}{2}}^{\mu,\pm}}}.$
The PP property of the updated cell averages in scheme (59) is shown as follows.